### v1

# DM-MPPI: Datamodel for Efficient and Safe Model Path Integral Control

Abstract
We extend the Datamodels framework from machine learning to Model Predictive Path Integral (MPPI) control. While datamodels estimate sample influence via regression on fixed datasets, our approach learns to predict influence from sample cost features—enabling real-time estimation for new samples without online regression. The predictor is trained offline on influence coefficients computed via the datamodel framework across diverse MPPI instances, then deployed for efficient sample pruning and adaptive constraint handling. A single learned model addresses both efficiency and safety: low-influence samples are pruned to reduce computation, while monitoring constraint-violation influence enables adaptive penalty tuning. Experiments on path tracking with obstacle avoidance demonstrate $5\times$ sample reduction while maintaining performance and improving constraint satisfaction.
I Introduction
Model Predictive Path Integral (MPPI) control has emerged as a powerful framework for sampling-based optimal control in nonlinear systems [1, 2]. By generating trajectory samples and computing importance-weighted averages, MPPI avoids linearization while achieving real-time performance through parallelization [3]. However, two fundamental challenges remain. First, sample efficiency: many samples may contribute little to the final control, wasting computation. Second, constraint handling: soft constraint penalties are the standard approach [18], but understanding how constraint-violating samples influence the control output remains opaque.
The Datamodels framework from machine learning [8] offers a principled approach to understanding how individual data points affect model outputs. We adapt this framework to MPPI by learning to predict trajectory cost from sample inclusion. The key insight is simple: by predicting total cost, which includes constraint violation penalties, a single datamodel captures both task performance and constraint-relevant influence. The relative importance of constraints can be adjusted through the penalty coefficient without requiring separate models.
I-A Related Work
MPPI and Path Integral Control. Path integral methods derive optimal control laws through exponential weighting of sampled trajectories [4]. Williams et al. [1, 2] developed MPPI for real-time control with GPU parallelization, demonstrating aggressive autonomous driving [3]. Recent variants address robustness [5], smoothness [6], and sample efficiency through unscented guidance [7]. Our work complements these advances by providing interpretable influence analysis.
Data Attribution in Machine Learning. Understanding how training data affects model predictions is a fundamental problem in ML. Influence functions [9] adapt classical robust statistics to deep learning, while Data Shapley [10] applies game-theoretic valuation. TracIn [11] provides efficient gradient-based attribution. The Datamodels framework [8] learns linear predictors of model outputs from training set composition, achieving high accuracy on image classification. TRAK [12] scales attribution to large models via random projections. We are the first to apply datamodels to sampling-based control.
Constraint Handling in MPPI. Standard MPPI handles constraints through soft penalties in the cost function, but this approach cannot guarantee constraint satisfaction and requires careful tuning of penalty weights. To address this limitation, several MPPI-specific methods have been developed. Risk-Aware MPPI (RA-MPPI) [13] replaces the expected cost with Conditional Value-at-Risk (CVaR), steering optimization toward the worst-case tail of the rollout distribution to improve safety in autonomous racing. Shield-MPPI [14] integrates Control Barrier Functions (CBFs) into MPPI through a dual-layer structure that combines cost penalization with a gradient-based local repair step, significantly reducing constraint violations while maintaining computational efficiency on CPUs. For hard constraint satisfaction, Constrained Covariance Steering MPPI (CCS-MPPI) [15] enforces state and control constraints through a tube-based formulation, though it operates probabilistically. Gandhi et al. [22] develop safe importance sampling that biases trajectory generation toward constraint-satisfying regions while preserving theoretical guarantees. Our datamodel-based approach complements these methods by diagnosing how constraint-violating samples influence the control output, enabling principled tuning of penalty parameters rather than enforcing constraints directly.
Learning for Control. Deep learning increasingly intersects with control theory [23, 24]. Data-driven methods can directly synthesize controllers from data [25]. Neural certificates [26] learn Lyapunov and barrier functions alongside policies. Interpretability of learned policies is an emerging concern [27]. Our work contributes to this intersection by bringing data attribution tools to sampling-based control.
I-B Contributions
This paper extends the Datamodels framework to MPPI by learning to predict sample influence from cost features rather than computing influence for fixed datasets. Our key contribution is an offline-online architecture: ground-truth influence coefficients are computed via datamodel regression during training, then predicted via a neural network at runtime—enabling real-time estimation without online regression. A single learned predictor addresses both sample efficiency through pruning and constraint handling through adaptive penalty tuning.
The remainder of this paper is organized as follows. Section II provides background on MPPI control and the Datamodels framework. Section III presents the proposed MPPI-Datamodel framework, detailing the offline training procedure for learning the influence predictor and the online inference phase for sample pruning and adaptive constraint handling. Section IV evaluates the framework on a path tracking task with obstacle avoidance, demonstrating sample efficiency gains and improved safety margins. Section V discusses limitations and directions for future work, and Section VI concludes the paper.
II Preliminaries
II-A Model Predictive Path Integral Control
Consider a nonlinear stochastic system with control-affine dynamics
$dx=f(x)dt+G(x)(u+\epsilon)dt$ (1)
where $x\in\mathbb{R}^{n}$ is the state, $u\in\mathbb{R}^{m}$ is the control input, $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is the drift dynamics, $G:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ is the control matrix, and $\epsilon\sim\mathcal{N}(0,\Sigma)$ represents control noise.
The total trajectory cost consists of multiple components
$S(\tau)=S_{\text{goal}}(\tau)+S_{\text{control}}(\tau)+\rho\cdot S_{\text{violation}}(\tau)$ (2)
where the goal cost captures task objectives $S_{\text{goal}}(\tau)=\phi(x_{T})+\sum_{t=0}^{T-1}q(x_{t})$, the control cost penalizes actuation $S_{\text{control}}(\tau)=\sum_{t=0}^{T-1}\frac{1}{2}u_{t}^{\top}Ru_{t}$, and the violation cost captures soft constraint penalties $S_{\text{violation}}(\tau)=\sum_{t=0}^{T-1}c(x_{t})$ with $\rho>0$ controlling constraint importance.
Given $K$ sampled trajectories $\{\tau_{k}\}_{k=1}^{K}$, each generated by applying perturbed controls $u_{t}^{k}=\bar{u}_{t}+\epsilon_{t}^{k}$, the optimal control update is
$u_{t}^{*}=\bar{u}_{t}+\sum_{k=1}^{K}w_{k}\epsilon_{t}^{k}$ (3)
where the importance weights follow the Gibbs distribution
$w_{k}=\frac{\exp(-S(\tau_{k})/\lambda)}{\sum_{j=1}^{K}\exp(-S(\tau_{j})/\lambda)}.$ (4)
The temperature $\lambda>0$ governs weight concentration [2].
II-B The Datamodels Framework
The Datamodels framework [8] provides a principled approach to understanding how individual data points influence model outputs. The core idea is to learn a simple predictor that maps training set composition to model behavior, enabling post-hoc attribution of predictions to specific training examples.
For a dataset $\mathcal{S}=\{z_{1},\ldots,z_{n}\}$ and a model output of interest $f(x;S^{\prime})$ trained on subset $S^{\prime}\subseteq\mathcal{S}$, a linear datamodel takes the form
$g_{\theta}(\mathbf{1}_{S^{\prime}})=\theta^{\top}\mathbf{1}_{S^{\prime}}+\theta_{0}$ (5)
where $\mathbf{1}_{S^{\prime}}\in\{0,1\}^{n}$ is a binary vector indicating which data points are included in subset $S^{\prime}$. The coefficient $\theta_{j}$ quantifies the marginal influence of data point $z_{j}$ on the output: positive $\theta_{j}$ indicates that including $z_{j}$ increases the predicted output, while negative $\theta_{j}$ indicates a decreasing effect.
Training proceeds by constructing many random subsets $\{S^{\prime}_{i}\}_{i=1}^{M}$, each formed by including each data point independently with probability $\alpha\in(0,1)$. For each subset, the true model output $f(x;S^{\prime}_{i})$ is computed. The datamodel parameters are then learned via LASSO regression [12]:
$\theta^{*}=\operatorname*{arg\,min}_{\theta}\frac{1}{M}\sum_{i=1}^{M}\left(g_{\theta}(\mathbf{1}_{S^{\prime}_{i}})-f(x;S^{\prime}_{i})\right)^{2}+\mu\|\theta\|_{1}$ (6)
The $\ell_{1}$ penalty encourages sparsity, reflecting the intuition that only a subset of data points meaningfully influence any given prediction.
Unlike influence functions [9], which rely on local approximations around the full training set, datamodels directly learn global input-output relationships across the space of possible training subsets.
II-C Why Linear Models Suit MPPI
The weighted cost can be written as
$\bar{S}(\mathbf{b})=\frac{\sum_{k}b_{k}e^{-S_{k}/\lambda}S_{k}}{\sum_{j}b_{j}e^{-S_{j}/\lambda}}=\frac{N(\mathbf{b})}{D(\mathbf{b})}$ (7)
where both numerator $N(\mathbf{b})$ and denominator $D(\mathbf{b})$ are linear in the inclusion vector $\mathbf{b}$, but their ratio is not. However, when subsets are formed by including each sample independently with probability $\alpha$, the denominator concentrates around its expectation for sufficiently large $\alpha K$:
$D(\mathbf{b})\approx\mathbb{E}[D]=\alpha\sum_{k}e^{-S_{k}/\lambda}.$ (8)
Substituting this approximation yields
$\bar{S}(\mathbf{b})\approx\frac{1}{\mathbb{E}[D]}\sum_{k}b_{k}e^{-S_{k}/\lambda}S_{k}=\sum_{k}\theta_{k}b_{k}$ (9)
where $\theta_{k}\propto e^{-S_{k}/\lambda}S_{k}$. Thus, when subset size is large enough for the denominator to concentrate, the weighted cost becomes approximately linear in $\mathbf{b}$. This approximation degrades when $\alpha K$ is small or when a few samples dominate the weights (low $\lambda$, high cost variance), causing $D(\mathbf{b})$ to fluctuate significantly across subsets. Nonetheless, even when these conditions are not fully met, the linear datamodel can still serve as a practical approximation worth investigating.
III Method: MPPI-Datamodel Framework
[…]
Figure 1: MPPI-Datamodel workflow. Offline phase (green): collect MPPI instances, fit datamodels via LASSO regression, and train influence predictor $h_{\phi}$. Online phase: predict influence using $h_{\phi^{*}}$ with cost features and instance statistics, prune low-influence samples for efficiency, and adapt penalty $\rho$ based on violation influence ratio for safety.
III-A Problem Formulation
Consider an MPPI instance with $K$ sampled trajectories $\mathcal{T}=\{\tau_{1},\ldots,\tau_{K}\}$. Each trajectory $\tau_{k}$ incurs a total cost $C_{k}$ composed of goal cost, control cost, and constraint violation cost:
$C_{k}=C_{\text{goal},k}+C_{\text{ctrl},k}+\rho\cdot C_{\text{viol},k}$ (10)
where $\rho>0$ is the soft constraint penalty coefficient. For a subset $\mathcal{T}^{\prime}\subseteq\mathcal{T}$, the subset-restricted weighted cost is
$\bar{C}(\mathcal{T}^{\prime})=\sum_{k:\tau_{k}\in\mathcal{T}^{\prime}}w_{k}(\mathcal{T}^{\prime})\,C_{k}$ (11)
with renormalized importance weights
$w_{k}(\mathcal{T}^{\prime})=\frac{\exp(-C_{k}/\lambda)}{\sum_{j:\tau_{j}\in\mathcal{T}^{\prime}}\exp(-C_{j}/\lambda)}$ (12)
To capture the relationship between sample inclusion and weighted cost, we define the binary inclusion vector $\mathbf{b}\in\{0,1\}^{K}$ where $b_{k}=1$ if trajectory $\tau_{k}$ is included. The MPPI-Datamodel predicts weighted cost from sample membership via the linear model
$g_{\theta}(\mathbf{b})=\theta^{\top}\mathbf{b}+\theta_{0}\approx\bar{C}(\mathcal{T}^{\prime})$ (13)
where the coefficient $\theta_{k}$ represents the marginal influence of sample $\tau_{k}$ on the weighted cost. Since the MPPI control update $u^{*}=\sum_{k}w_{k}u_{k}$ uses the same importance weights $w_{k}$, samples that significantly affect the weighted cost also significantly affect the control output. This justifies using influence on weighted cost as a proxy for influence on the control decision.
III-B Offline Training Phase
We adopt an offline training paradigm to avoid repeated regression during deployment. The key insight from the datamodeling framework [8] is that by training on many random subsets, we can recover individual sample influences even though we never isolate single samples. The offline procedure consists of three steps.
Step 1: Data Collection. We collect $N$ MPPI instances across representative states sampled from the expected operating distribution. For each instance $i$, we draw $K$ trajectories using the nominal sampling distribution and record the cost components $(C_{\text{goal},k}^{(i)},C_{\text{ctrl},k}^{(i)},C_{\text{viol},k}^{(i)})$ for each trajectory $k$. Given a fixed penalty $\rho_{0}$, we compute total costs $C_{k}^{(i)}=C_{\text{goal},k}^{(i)}+C_{\text{ctrl},k}^{(i)}+\rho_{0}\cdot C_{\text{viol},k}^{(i)}$. To ensure the learned predictor generalizes, instances should cover diverse regions of the state space, including states near constraints.
Step 2: Datamodel Fitting. For each instance $i$, we learn the influence coefficients by regression on random subsets. We generate $M$ random subsets by including each trajectory independently with probability $\alpha\in(0,1)$. For each subset $j$ with inclusion mask $\mathbf{b}_{j}\in\{0,1\}^{K}$, we compute the subset-restricted weighted cost
$\bar{C}_{j}^{(i)}=\frac{\sum_{k=1}^{K}b_{j,k}\,w_{k}^{(i)}\,C_{k}^{(i)}}{\sum_{k=1}^{K}b_{j,k}\,w_{k}^{(i)}}$ (14)
where $w_{k}^{(i)}=\exp(-C_{k}^{(i)}/\lambda)$ are the unnormalized importance weights. The influence coefficients are obtained by solving the LASSO regression problem
$\theta^{(i)*}=\operatorname*{arg\,min}_{\theta,\theta_{0}}\frac{1}{M}\sum_{j=1}^{M}\left(\theta^{\top}\mathbf{b}_{j}+\theta_{0}-\bar{C}_{j}^{(i)}\right)^{2}+\mu\|\theta\|_{1}$ (15)
where $\theta_{0}$ is an intercept term excluded from regularization and the $\ell_{1}$ penalty encourages sparsity.
The key to recovering individual influences is the randomness of the subsets. For any two trajectories $k\neq k^{\prime}$, their inclusion events are independent across subsets:
$\mathbb{E}[b_{j,k}\cdot b_{j,k^{\prime}}]=\mathbb{E}[b_{j,k}]\cdot\mathbb{E}[b_{j,k^{\prime}}]=\alpha^{2}$ (16)
This independence decorrelates the contributions of different samples, allowing the regression to disentangle individual effects from aggregate subset effects. With sufficient subsets $M$, the LASSO solution $\theta_{k}^{(i)*}$ accurately estimates the marginal influence of each sample $k$.
Step 3: Influence Predictor Training. We train a predictor $h_{\phi}$ that generalizes influence estimation across instances. A critical observation is that the influence coefficient $\theta_{k}^{(i)*}$ depends not only on sample $k$’s own cost, but also on the cost distribution of other samples in instance $i$. Intuitively, a trajectory with cost $C_{k}=10$ has high influence if all other trajectories have costs above $100$, but low influence if other trajectories have similar costs around $10$.
To capture this context dependence, we augment the input features with summary statistics:
$\hat{\theta}_{k}=h_{\phi}\left(C_{k},C_{\text{viol},k},\bar{C}^{(i)},\sigma_{C}^{(i)}\right)$ (17)
where $\bar{C}^{(i)}=\frac{1}{K}\sum_{k=1}^{K}C_{k}^{(i)}$ is the mean cost and $\sigma_{C}^{(i)}=\text{std}(\{C_{k}^{(i)}\}_{k=1}^{K})$ is the standard deviation. These statistics provide context about where sample $k$ sits relative to other samples in the instance.
The predictor parameters are learned by minimizing prediction error across all collected instances:
$\phi^{*}=\operatorname*{arg\,min}_{\phi}\sum_{i=1}^{N}\sum_{k=1}^{K}\left(h_{\phi}\left(C_{k}^{(i)},C_{\text{viol},k}^{(i)},\bar{C}^{(i)},\sigma_{C}^{(i)}\right)-\theta_{k}^{(i)*}\right)^{2}$ (18)
In practice, $h_{\phi}$ can be a small neural network (e.g., two hidden layers with 64 units) or even a linear model if the relationship is sufficiently simple. Algorithm 1 summarizes the complete offline training procedure.
Algorithm 1 Offline: Influence Predictor Training
0: Number of instances NN, samples per instance $K$, subsets per instance $M$, inclusion probability $\alpha$, regularization $\mu$, initial penalty $\rho_{0}$
0: Trained predictor $h_{\phi^{*}}$
1: Initialize dataset $\mathcal{D}\leftarrow\emptyset$
2:for $i=1$ to NN do
3:  Sample state $x_{0}^{(i)}$ from operating distribution
4:  Collect $K$ trajectories, compute cost components $\{C_{\text{goal},k}^{(i)},C_{\text{ctrl},k}^{(i)},C_{\text{viol},k}^{(i)}\}_{k=1}^{K}$
5:  Compute total costs: $C_{k}^{(i)}\leftarrow C_{\text{goal},k}^{(i)}+C_{\text{ctrl},k}^{(i)}+\rho_{0}\cdot C_{\text{viol},k}^{(i)}$
6:  Compute importance weights: $w_{k}^{(i)}\leftarrow\exp(-C_{k}^{(i)}/\lambda)$
7:  Compute instance statistics: $\bar{C}^{(i)}\leftarrow\frac{1}{K}\sum_{k}C_{k}^{(i)}$,  $\sigma_{C}^{(i)}\leftarrow\text{std}(\{C_{k}^{(i)}\})$
8:  for $j=1$ to $M$ do
9:   Sample inclusion mask: $b_{j,k}\sim\text{Bernoulli}(\alpha)$ for all $k$
10:   Compute weighted cost: $\bar{C}_{j}^{(i)}\leftarrow\frac{\sum_{k}b_{j,k}\,w_{k}^{(i)}\,C_{k}^{(i)}}{\sum_{k}b_{j,k}\,w_{k}^{(i)}}$
11:  end for
12:  Solve LASSO (15) to obtain $\theta^{(i)*}\in\mathbb{R}^{K}$
13:  for $k=1$ to $K$ do
14:   $\mathcal{D}\leftarrow\mathcal{D}\cup\{(C_{k}^{(i)},C_{\text{viol},k}^{(i)},\bar{C}^{(i)},\sigma_{C}^{(i)},\theta_{k}^{(i)*})\}$
15:  end for
16:end for
17: Train $h_{\phi^{*}}$ on $\mathcal{D}$ by minimizing (18)
18:return $h_{\phi^{*}}$
III-C Online Inference Phase
At runtime, the learned predictor enables efficient influence estimation without solving any regression problems. The predictor uses the decomposed costs $(C_{k},C_{\text{viol},k})$ along with instance statistics as input features. Importantly, the actual MPPI control computation still operates on the total cost with the current penalty $\rho$, which can be adjusted online without retraining the predictor.
Given a new set of $K$ sampled trajectories, we first compute the cost components $(C_{\text{goal},k},C_{\text{ctrl},k},C_{\text{viol},k})$ and total costs $C_{k}=C_{\text{goal},k}+C_{\text{ctrl},k}+\rho\cdot C_{\text{viol},k}$. We then compute the instance statistics
$\bar{C}=\frac{1}{K}\sum_{k=1}^{K}C_{k},\qquad\sigma_{C}=\text{std}(\{C_{k}\}_{k=1}^{K})$ (19)
and obtain the predicted influence coefficients as
$\hat{\theta}_{k}=h_{\phi^{*}}(C_{k},C_{\text{viol},k},\bar{C},\sigma_{C})$ (20)
These predictions identify which samples contribute meaningfully to the control output.
Sample Pruning for Efficiency. We retain only those samples whose predicted influence magnitude exceeds a threshold $\tau$:
$\mathcal{K}_{\text{keep}}=\{k:|\hat{\theta}_{k}|\geq\tau\}$ (21)
This pruning step reduces computational cost while preserving the samples that matter most. The MPPI control update is then computed using only the retained samples:
$u^{*}=\sum_{k\in\mathcal{K}_{\text{keep}}}\tilde{w}_{k}\,u_{k},\qquad\tilde{w}_{k}=\frac{\exp(-C_{k}/\lambda)}{\sum_{j\in\mathcal{K}_{\text{keep}}}\exp(-C_{j}/\lambda)}$ (22)
where the importance weights are renormalized over $\mathcal{K}_{\text{keep}}$.
Adaptive Constraint Penalty for Safety. To adapt the constraint penalty online, we monitor the influence contributed by constraint-violating samples. We define the violation influence ratio as
$r_{\text{viol}}=\frac{\sum_{k\in\mathcal{K}_{\text{viol}}}|\hat{\theta}_{k}|}{\sum_{k=1}^{K}|\hat{\theta}_{k}|}$ (23)
where $\mathcal{K}_{\text{viol}}=\{k:C_{\text{viol},k}>0\}$ denotes the set of constraint-violating samples. This ratio quantifies how much of the total influence comes from samples that violate constraints.
When $r_{\text{viol}}$ is high, constraint-violating trajectories disproportionately affect the control output, suggesting the current penalty $\rho$ is insufficient to discourage unsafe behaviors. We update the penalty coefficient according to
$\rho_{t+1}=\rho_{t}+\eta\cdot(r_{\text{viol}}-r_{\text{target}})$ (24)
where $\eta>0$ is a step size and $r_{\text{target}}\in(0,1)$ is the desired maximum violation influence ratio. This adaptation increases $\rho$ when violation influence exceeds the target and decreases it otherwise, maintaining constraint satisfaction while avoiding overly conservative behavior. Algorithm 2 presents the complete online procedure.
Algorithm 2 Online: Datamodel-Based MPPI
0: Trained predictor $h_{\phi^{*}}$, initial state $x_{0}$, nominal control $\bar{U}$, initial penalty $\rho$, pruning threshold $\tau$, target violation ratio $r_{\text{target}}$, adaptation step size $\eta$, temperature $\lambda$
1:for each control iteration do
2:  Sample $K$ trajectories from current state
3:  Compute cost components: $(C_{\text{goal},k},C_{\text{ctrl},k},C_{\text{viol},k})$ for all $k$
4:  Compute total costs: $C_{k}\leftarrow C_{\text{goal},k}+C_{\text{ctrl},k}+\rho\cdot C_{\text{viol},k}$
5:  Compute instance statistics: $\bar{C}\leftarrow\frac{1}{K}\sum_{k}C_{k}$,  $\sigma_{C}\leftarrow\text{std}(\{C_{k}\})$
6:  Predict influence: $\hat{\theta}_{k}\leftarrow h_{\phi^{*}}(C_{k},C_{\text{viol},k},\bar{C},\sigma_{C})$ for all $k$
7:  Identify kept samples: $\mathcal{K}_{\text{keep}}\leftarrow\{k:|\hat{\theta}_{k}|\geq\tau\}$
8:  Identify violating samples: $\mathcal{K}_{\text{viol}}\leftarrow\{k:C_{\text{viol},k}>0\}$
9:  Compute violation ratio: $r_{\text{viol}}\leftarrow\frac{\sum_{k\in\mathcal{K}_{\text{viol}}}|\hat{\theta}_{k}|}{\sum_{k=1}^{K}|\hat{\theta}_{k}|}$
10:  Update penalty: $\rho\leftarrow\rho+\eta\cdot(r_{\text{viol}}-r_{\text{target}})$
11:  Compute weights: $\tilde{w}_{k}\leftarrow\frac{\exp(-C_{k}/\lambda)}{\sum_{j\in\mathcal{K}_{\text{keep}}}\exp(-C_{j}/\lambda)}$ for $k\in\mathcal{K}_{\text{keep}}$
12:  Compute optimal control: $u^{*}\leftarrow\sum_{k\in\mathcal{K}_{\text{keep}}}\tilde{w}_{k}\,u_{k}$
13:  Execute first control $u_{0}^{*}$, shift nominal sequence $\bar{U}$
14:end for
Remark 1 (Unified Influence for Efficiency and Safety).
Because the predictor takes both total cost and violation cost as input, a single learned model captures influence relevant to both sample efficiency and constraint handling. The pruning mechanism improves efficiency by discarding low-influence samples regardless of their constraint status, while the adaptive $\rho$ mechanism improves safety by strengthening penalties when constraint-violating samples remain influential. These two mechanisms operate complementarily: pruning reduces computation without sacrificing control quality, while penalty adaptation ensures the kept samples lead to safe behaviors.
Remark 2 (Computational Complexity).
The offline phase requires $O(N\cdot M\cdot K)$ subset evaluations and $N$ LASSO solves, but is performed only once. The online phase requires only $O(K)$ forward passes through $h_{\phi^{*}}$ per control iteration, making influence prediction negligible compared to trajectory sampling and cost evaluation. The pruning step further reduces the cost of the weighted average computation from $O(K)$ to $O(|\mathcal{K}_{\text{keep}}|)$.
IV Experiments
We evaluate the proposed framework on a path tracking task with obstacle avoidance, demonstrating unified benefits for sample efficiency and constraint-aware control.
IV-A Setup
We consider a vehicle governed by the kinematic bicycle model with state $x=[p_{x},p_{y},\psi,v]^{\top}$ (position, heading, velocity) and control $u=[\delta,a]^{\top}$ (steering angle, acceleration). The vehicle tracks an oval reference path while avoiding circular obstacles. The goal cost penalizes deviation from the reference path, while the violation cost applies a quadratic penalty when the vehicle enters the obstacle regions.
The MPPI controller uses a planning horizon of $T=20$ steps at $dt=0.1$ s and temperature $\lambda=100$. Control noise is sampled from $\mathcal{N}(0,\Sigma)$ with $\Sigma=\text{diag}(0.1,0.5)$ for steering and acceleration respectively.
For the influence predictor $h_{\phi}$, we use a multilayer perceptron (MLP) with two hidden layers of 64 units each and ReLU activations. The network takes as input the normalized costs $(S_{k},S_{\text{viol},k},\bar{S},\sigma_{S})$ and outputs the predicted influence coefficient $\hat{\theta}_{k}$. Training uses the Adam optimizer with learning rate $10^{-3}$ for 1000 epochs. The datamodel parameters are set as: inclusion probability $\alpha=0.5$, number of subsets $M=50$, and LASSO regularization $\mu=0.01$. We collect $N=200$ MPPI instances from randomly sampled initial states for offline training.
IV-B Sample Efficiency Analysis
Figure 2: Sample efficiency comparison between DM-MPPI and standard MPPI across varying sample sizes $K\in\{50,100,150,\ldots,500\}$.
To determine the minimum number of samples required for effective control, we evaluate both methods across sample sizes $K\in\{50,100,150,\ldots,500\}$. Figure 2 presents the results. Standard MPPI shows a steep performance degradation below $K=200$, with position RMSE increasing from 0.41 m at $K=500$ to 1.30 m at $K=50$. In contrast, DM-MPPI maintains stable performance across a wider range: at $K=100$, it achieves position RMSE of 0.44 m and $r_{\text{viol}}=0.11$, comparable to standard MPPI at $K=500$ (0.41 m and 0.09 respectively).
However, the datamodel approach has a minimum sample requirement. At $K=50$, DM-MPPI performance degrades noticeably (position RMSE increases to 0.72 m, $r_{\text{viol}}$ rises to 0.35), suggesting insufficient samples for reliable influence estimation. This aligns with the effective influence size $K_{\text{eff}}^{\theta}\approx 28$ observed in our analysis—when the total sample count approaches this threshold, the datamodel cannot effectively distinguish high-influence samples from noise.
Based on this analysis, we select $K=100$ for DM-MPPI in subsequent experiments, achieving $5\times$ sample reduction compared to standard MPPI ($K=500$) while maintaining equivalent control quality.
IV-C Performance Comparison
Figure 3: Comparison of sampled different trajectories.
TABLE I: Performance Comparison
Metric Standard DM-Fixed DM-Adaptive
Samples used 500 100 100
Position RMSE (m) 0.41 0.44 0.43
Heading RMSE (rad) 0.031 0.033 0.032
Min. obstacle dist. (m) 1.24 1.31 1.42
$r_{\text{viol}}$ 0.091 0.082 0.071
Iteration time (ms) 85.4 18.2 18.5
We compare three configurations: (1) Standard MPPI with $K=500$ samples and fixed penalty $\rho=10^{10}$, (2) DM-MPPI with $K=100$ samples and fixed $\rho=10^{10}$, and (3) DM-MPPI with $K=100$ samples and adaptive $\rho$ using $r_{\text{target}}=0.05$ and step size $\eta=10^{9}$.
Table I summarizes the performance comparison. The datamodel approach enables $5\times$ sample reduction while maintaining tracking quality. Figure 3 visualizes the sampled trajectories at a representative timestep; despite using only 100 samples, the datamodel variant produces a control output comparable to standard MPPI with 500 samples. Analysis of the learned influence coefficients reveals an effective influence size of $K_{\text{eff}}^{\theta}\approx 28$, confirming that roughly one-quarter of the samples carry meaningful influence while the remainder can be safely pruned. The adaptive variant achieves the largest minimum obstacle distance by dynamically adjusting $\rho$ based on the violation influence ratio $r_{\text{viol}}$.
V Limitations and Future Work
Soft Constraint Guarantees. The proposed framework relies on soft constraint penalties, which cannot guarantee hard constraint satisfaction. While the adaptive $\rho$ mechanism reduces the influence of constraint-violating samples, it cannot ensure that the resulting control will never violate constraints. Safety-critical applications may require integration with formal methods such as Control Barrier Functions [20, 21] or tube-based robust MPC [17]. Future work could integrate the datamodel framework with CBFs, using predicted influence to inform barrier parameters.
Limited Experimental Scope. The current evaluation considers a single path tracking task with obstacle avoidance using a kinematic bicycle model. While this serves as a representative benchmark, comprehensive validation requires evaluation across diverse domains including manipulation, legged locomotion, and higher-dimensional systems with more complex constraint structures. Future work will include comparisons against state-of-the-art sample-efficient MPPI variants [7, 16], evaluation under varying noise levels and constraint complexities, and deployment on real hardware platforms.
Fixed Predictor After Offline Training. The influence predictor is fixed after offline training, which limits adaptability to changing conditions. Online fine-tuning would allow the predictor to adapt to distribution shifts or novel scenarios, but introduces computational overhead that conflicts with real-time requirements. One promising direction is to maintain a small replay buffer of recent MPPI instances and perform incremental updates during idle cycles. Alternatively, meta-learning approaches could train the predictor to rapidly adapt to new conditions with minimal fine-tuning samples.
Generalization Across Cost Scales. The framework uses a fixed inclusion probability $\alpha$ during offline training, and the learned predictor generalizes well within the range of cost distributions encountered during training. However, performance may degrade when deployed in scenarios where cost statistics differ substantially from training. Normalizing costs by instance-level statistics partially addresses this issue, but does not fully resolve it when the underlying cost structure changes significantly. Future work could explore scale-invariant predictor architectures that operate on relative cost rankings rather than absolute values to enable broader generalization.
VI Conclusion
We presented a framework that extends the datamodeling paradigm from machine learning to sampling-based optimal control. While traditional datamodels estimate sample influence through regression on a fixed dataset, our approach trains an influence predictor that generalizes across control instances—enabling real-time influence estimation without solving regression problems online. By learning to map cost features to influence coefficients, the predictor identifies which samples contribute meaningfully to the control output, allowing efficient sample pruning and adaptive constraint handling through a single learned model.
The key insight is that influence coefficients, though computed via subset regression during offline training, show predictable structure as a function of sample costs and global statistics. This structure enables amortized inference: rather than re-running regression for each new MPPI instance, we predict influence directly from cost features in a single forward pass. Experiments on path tracking with obstacle avoidance demonstrate $5\times$ sample reduction while maintaining tracking performance and achieving improved safety margins through adaptive penalty tuning.

### v2

@@
-DM-MPPI: Datamodel for Efficient and Safe Model Path Integral Control
+DM-MPPI: Datamodel-Based Sample Pruning for Model Predictive Path Integral Control
 Abstract
-We extend the Datamodels framework from machine learning to Model Predictive Path Integral (MPPI) control.
-While datamodels estimate sample influence via regression on fixed datasets, our approach learns to predict influence from sample cost features—enabling real-time estimation for new samples without online regression.
-The predictor is trained offline on influence coefficients computed via the datamodel framework across diverse MPPI instances, then deployed for efficient sample pruning and adaptive constraint handling.
-A single learned model addresses both efficiency and safety: low-influence samples are pruned to reduce computation, while monitoring constraint-violation influence enables adaptive penalty tuning.
-Experiments on path tracking with obstacle avoidance demonstrate $5\times$ sample reduction while maintaining performance and improving constraint satisfaction.
+Sampling-based controllers like Model Predictive Path Integral (MPPI) generate hundreds of trajectory rollouts every control cycle, but many of these rollouts barely affect—or even hurt—the final control action.
+Figuring out which samples actually matter is, at its core, a data attribution problem: each rollout acts like a training example that pushes the controller’s output one way or another, depending on whether it is included.
+We make this connection concrete by adapting the Datamodels framework, which learns how much each sample matters through regression over random subsets, to MPPI.
+During an offline phase, we compute influence scores across random subsets of rollouts and train a small neural network to predict influence from cost features—allowing real-time pruning without any online regression.
+An adaptive constraint penalty also adjusts obstacle avoidance strength based on how often sampled trajectories violate constraints.
+We validate the method on a kinematic bicycle model driving through two obstacle-filled tracks.
 I Introduction
-Model Predictive Path Integral (MPPI) control has emerged as a powerful framework for sampling-based optimal control in nonlinear systems [1, 2].
-By generating trajectory samples and computing importance-weighted averages, MPPI avoids linearization while achieving real-time performance through parallelization [3].
-However, two fundamental challenges remain.
-First, sample efficiency: many samples may contribute little to the final control, wasting computation.
-Second, constraint handling: soft constraint penalties are the standard approach [18], but understanding how constraint-violating samples influence the control output remains opaque.
-The Datamodels framework from machine learning [8] offers a principled approach to understanding how individual data points affect model outputs.
-We adapt this framework to MPPI by learning to predict trajectory cost from sample inclusion.
-The key insight is simple: by predicting total cost, which includes constraint violation penalties, a single datamodel captures both task performance and constraint-relevant influence.
-The relative importance of constraints can be adjusted through the penalty coefficient without requiring separate models.
-I-A Related Work
-MPPI and Path Integral Control.
-Path integral methods derive optimal control laws through exponential weighting of sampled trajectories [4].
+Model Predictive Path Integral (MPPI) control has become a popular sampling-based approach for nonlinear systems [17, 16], achieving real-time performance by rolling out many trajectories in parallel and blending them through importance-weighted averaging [18].
+However, two key challenges remain: a large fraction of the sampled trajectories has little effect on the final control action, leading to wasted computation, and soft constraint penalties [23] make it unclear how constraint-violating samples shape the output.
+A natural question is whether we can identify which samples actually matter and focus the computation on those.
+We argue that this is naturally a data attribution problem — each rollout is like a training example that, when included or left out, shifts the control update in a different direction.
+The Datamodels framework [7] offers a clean solution by learning per-example influence coefficients through subset regression.
+We bring this idea into MPPI by moving the expensive regression to an offline stage and training a small neural network to predict influence at deployment time.
+Path integral methods recover optimal control laws by exponentially weighting sampled trajectories [14].
 Williams et al.
-[1, 2] developed MPPI for real-time control with GPU parallelization, demonstrating aggressive autonomous driving [3].
-Recent variants address robustness [5], smoothness [6], and sample efficiency through unscented guidance [7].
-Our work complements these advances by providing interpretable influence analysis.
-Data Attribution in Machine Learning.
-Understanding how training data affects model predictions is a fundamental problem in ML.
-Influence functions [9] adapt classical robust statistics to deep learning, while Data Shapley [10] applies game-theoretic valuation.
-TracIn [11] provides efficient gradient-based attribution.
-The Datamodels framework [8] learns linear predictors of model outputs from training set composition, achieving high accuracy on image classification.
-TRAK [12] scales attribution to large models via random projections.
-We are the first to apply datamodels to sampling-based control.
-Constraint Handling in MPPI.
-Standard MPPI handles constraints through soft penalties in the cost function, but this approach cannot guarantee constraint satisfaction and requires careful tuning of penalty weights.
-To address this limitation, several MPPI-specific methods have been developed.
-Risk-Aware MPPI (RA-MPPI) [13] replaces the expected cost with Conditional Value-at-Risk (CVaR), steering optimization toward the worst-case tail of the rollout distribution to improve safety in autonomous racing.
-Shield-MPPI [14] integrates Control Barrier Functions (CBFs) into MPPI through a dual-layer structure that combines cost penalization with a gradient-based local repair step, significantly reducing constraint violations while maintaining computational efficiency on CPUs.
-For hard constraint satisfaction, Constrained Covariance Steering MPPI (CCS-MPPI) [15] enforces state and control constraints through a tube-based formulation, though it operates probabilistically.
-Gandhi et al.
-[22] develop safe importance sampling that biases trajectory generation toward constraint-satisfying regions while preserving theoretical guarantees.
-Our datamodel-based approach complements these methods by diagnosing how constraint-violating samples influence the control output, enabling principled tuning of penalty parameters rather than enforcing constraints directly.
-Learning for Control.
-Deep learning increasingly intersects with control theory [23, 24].
-Data-driven methods can directly synthesize controllers from data [25].
-Neural certificates [26] learn Lyapunov and barrier functions alongside policies.
-Interpretability of learned policies is an emerging concern [27].
-Our work contributes to this intersection by bringing data attribution tools to sampling-based control.
-I-B Contributions
-This paper extends the Datamodels framework to MPPI by learning to predict sample influence from cost features rather than computing influence for fixed datasets.
-Our key contribution is an offline-online architecture: ground-truth influence coefficients are computed via datamodel regression during training, then predicted via a neural network at runtime—enabling real-time estimation without online regression.
-A single learned predictor addresses both sample efficiency through pruning and constraint handling through adaptive penalty tuning.
-The remainder of this paper is organized as follows.
-Section II provides background on MPPI control and the Datamodels framework.
-Section III presents the proposed MPPI-Datamodel framework, detailing the offline training procedure for learning the influence predictor and the online inference phase for sample pruning and adaptive constraint handling.
-Section IV evaluates the framework on a path tracking task with obstacle avoidance, demonstrating sample efficiency gains and improved safety margins.
-Section V discusses limitations and directions for future work, and Section VI concludes the paper.
+[17, 16] introduced MPPI for real-time GPU-parallel control and showed its effectiveness in aggressive autonomous driving [18].
+Follow-up work has since improved robustness [5], smoothness [8], and sample efficiency [11].
+On the constraint side, researchers have explored CVaR-based risk awareness [22], Control Barrier Functions [21], tube-based formulations [20], and safe importance sampling [4].
+Deep learning intersects with control through data-driven synthesis [3, 10, 19] and neural certificates [2], with interpretability drawing growing attention [15].
+In the machine learning literature, data attribution has been studied through influence functions [9], Data Shapley [6], TracIn [13], the Datamodels framework [7], and TRAK [12].
+To our knowledge, no prior work has applied data attribution to sampling-based control.
+Our contributions are threefold.
+First, we recast MPPI sample selection as a data attribution problem and adapt the Datamodels framework through an offline–online setup: influence coefficients are fitted via subset regression during training and predicted by a neural network at runtime.
+Second, we compare nine methods on a kinematic bicycle model across two track geometries with obstacles.
+Keeping only $20\%$ of samples through informed pruning substantially lowers tracking error compared to unpruned MPPI, while random pruning is worse than even the smallest sample budget.
+Adding more samples shows diminishing returns, but pruning down to a few well-chosen samples consistently beats the full unpruned budget.
+Third, we show that the learned influence predictor with adaptive mechanisms (DM-adapted) achieves the fewest boundary violations across both tracks, significantly outperforming heuristic pruning methods that tend to keep trajectories running tight along obstacles—whereas influence-aware selection naturally avoids putting too much weight on samples near the constraint boundary.
+The rest of this paper is organized as follows.
+Section II covers background on MPPI and the Datamodels framework.
+Section III presents the proposed framework.
+Section IV evaluates against baselines on two track geometries.
+Section V concludes.
 II Preliminaries
 II-A Model Predictive Path Integral Control
-Consider a nonlinear stochastic system with control-affine dynamics
-$dx=f(x)dt+G(x)(u+\epsilon)dt$ (1)
-where $x\in\mathbb{R}^{n}$ is the state, $u\in\mathbb{R}^{m}$ is the control input, $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is the drift dynamics, $G:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ is the control matrix, and $\epsilon\sim\mathcal{N}(0,\Sigma)$ represents control noise.
+Consider a discrete-time nonlinear system with a control-affine structure.
+$x_{t+1}=F(x_{t},u_{t}+\epsilon_{t})$ (1)
+where $x_{t}\in\mathbb{R}^{n}$ is the state, $u_{t}\in\mathbb{R}^{m}$ is the control input, $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ is the dynamics map, and $\epsilon_{t}\sim\mathcal{N}(0,\Sigma)$ represents zero-mean Gaussian control noise with covariance $\Sigma\in\mathbb{R}^{m\times m}$.
 The total trajectory cost consists of multiple components
 $S(\tau)=S_{\text{goal}}(\tau)+S_{\text{control}}(\tau)+\rho\cdot S_{\text{violation}}(\tau)$ (2)
-where the goal cost captures task objectives $S_{\text{goal}}(\tau)=\phi(x_{T})+\sum_{t=0}^{T-1}q(x_{t})$, the control cost penalizes actuation $S_{\text{control}}(\tau)=\sum_{t=0}^{T-1}\frac{1}{2}u_{t}^{\top}Ru_{t}$, and the violation cost captures soft constraint penalties $S_{\text{violation}}(\tau)=\sum_{t=0}^{T-1}c(x_{t})$ with $\rho>0$ controlling constraint importance.
+where the goal cost captures task objectives $S_{\text{goal}}(\tau)=\phi(x_{T})+\sum_{t=0}^{T-1}q(x_{t})$, the control cost follows the standard MPPI formulation $S_{\text{control}}(\tau_{k})=\lambda\sum_{t=0}^{T-1}\bar{u}_{t}^{\top}R\,\epsilon_{t,k}$ (a cross-term between the nominal control and the perturbation [16]), and the violation cost captures soft constraint penalties $S_{\text{violation}}(\tau)=\sum_{t=0}^{T-1}c(x_{t})$ with $\rho>0$ controlling constraint importance.
 Given $K$ sampled trajectories $\{\tau_{k}\}_{k=1}^{K}$, each generated by applying perturbed controls $u_{t}^{k}=\bar{u}_{t}+\epsilon_{t}^{k}$, the optimal control update is
@@
 $g_{\theta}(\mathbf{1}_{S^{\prime}})=\theta^{\top}\mathbf{1}_{S^{\prime}}+\theta_{0}$ (5)
-where $\mathbf{1}_{S^{\prime}}\in\{0,1\}^{n}$ is a binary vector indicating which data points are included in subset $S^{\prime}$.
-The coefficient $\theta_{j}$ quantifies the marginal influence of data point $z_{j}$ on the output: positive $\theta_{j}$ indicates that including $z_{j}$ increases the predicted output, while negative $\theta_{j}$ indicates a decreasing effect.
-Training proceeds by constructing many random subsets $\{S^{\prime}_{i}\}_{i=1}^{M}$, each formed by including each data point independently with probability $\alpha\in(0,1)$.
+where $\mathbf{1}_{S^{\prime}}\in\{0,1\}^{n}$ is a binary vector indicating which data points are included in the subset $S^{\prime}$.
+The coefficient $\theta_{j}$ quantifies the marginal influence of data point $z_{j}$ on the output: a positive $\theta_{j}$ indicates that including $z_{j}$ increases the predicted output, while a negative $\theta_{j}$ indicates a decreasing effect.
+Training proceeds by constructing many random subsets $\{S^{\prime}_{i}\}_{i=1}^{M}$, each formed by including each data point independently with a probability $\alpha\in(0,1)$.
 For each subset, the true model output $f(x;S^{\prime}_{i})$ is computed.
@@
 Unlike influence functions [9], which rely on local approximations around the full training set, datamodels directly learn global input-output relationships across the space of possible training subsets.
-II-C Why Linear Models Suit MPPI
-The weighted cost can be written as
-$\bar{S}(\mathbf{b})=\frac{\sum_{k}b_{k}e^{-S_{k}/\lambda}S_{k}}{\sum_{j}b_{j}e^{-S_{j}/\lambda}}=\frac{N(\mathbf{b})}{D(\mathbf{b})}$ (7)
-where both numerator $N(\mathbf{b})$ and denominator $D(\mathbf{b})$ are linear in the inclusion vector $\mathbf{b}$, but their ratio is not.
-However, when subsets are formed by including each sample independently with probability $\alpha$, the denominator concentrates around its expectation for sufficiently large $\alpha K$:
-$D(\mathbf{b})\approx\mathbb{E}[D]=\alpha\sum_{k}e^{-S_{k}/\lambda}.$ (8)
-Substituting this approximation yields
-$\bar{S}(\mathbf{b})\approx\frac{1}{\mathbb{E}[D]}\sum_{k}b_{k}e^{-S_{k}/\lambda}S_{k}=\sum_{k}\theta_{k}b_{k}$ (9)
-where $\theta_{k}\propto e^{-S_{k}/\lambda}S_{k}$.
-Thus, when subset size is large enough for the denominator to concentrate, the weighted cost becomes approximately linear in $\mathbf{b}$.
-This approximation degrades when $\alpha K$ is small or when a few samples dominate the weights (low $\lambda$, high cost variance), causing $D(\mathbf{b})$ to fluctuate significantly across subsets.
-Nonetheless, even when these conditions are not fully met, the linear datamodel can still serve as a practical approximation worth investigating.
+II-C Motivation for Linear Influence Models in MPPI
+The MPPI control update (3) is a weighted sum of perturbations, where inclusion or exclusion of a sample changes both the numerator and the normalizing denominator of the weights.
+The control update is therefore not linear in the inclusion vector $\mathbf{b}\in\{0,1\}^{K}$, and there is no formal guarantee that the LASSO regression will recover accurate influence coefficients.
+Nonetheless, empirically we find that the linear datamodel achieves $R^{2}\approx 0.6$ when predicting the first-step control update from sample inclusion (Section IV).
+We treat this as a practical heuristic: the moderate $R^{2}$ indicates that a linear model captures some structure in sample influence, and the MLP predictor trained on these coefficients inherits both the signal and the approximation error.
+Whether the resulting pruning criterion is better than simple cost-based heuristics is an empirical question that we investigate in Section IV.
 III Method: MPPI-Datamodel Framework
-[…]
-Figure 1: MPPI-Datamodel workflow.
-Offline phase (green): collect MPPI instances, fit datamodels via LASSO regression, and train influence predictor $h_{\phi}$.
-Online phase: predict influence using $h_{\phi^{*}}$ with cost features and instance statistics, prune low-influence samples for efficiency, and adapt penalty $\rho$ based on violation influence ratio for safety.
+Figure 1: DM-MPPI framework.
 III-A Problem Formulation
@@
 $C_{k}=C_{\text{goal},k}+C_{\text{ctrl},k}+\rho\cdot C_{\text{viol},k}$ (7)
-where $\rho>0$ is the soft constraint penalty coefficient. For a subset $\mathcal{T}^{\prime}\subseteq\mathcal{T}$, the subset-restricted weighted cost is
-$\bar{C}(\mathcal{T}^{\prime})=\sum_{k:\tau_{k}\in\mathcal{T}^{\prime}}w_{k}(\mathcal{T}^{\prime})\,C_{k}$ (11)
-with renormalized importance weights
+where $\rho>0$ is the soft constraint penalty coefficient.
+For a subset $\mathcal{T}^{\prime}\subseteq\mathcal{T}$, the subset-restricted control update is computed by renormalizing importance weights over $\mathcal{T}^{\prime}$:
 $w_{k}(\mathcal{T}^{\prime})=\frac{\exp(-C_{k}/\lambda)}{\sum_{j:\tau_{j}\in\mathcal{T}^{\prime}}\exp(-C_{j}/\lambda)}$ (8)
-To capture the relationship between sample inclusion and weighted cost, we define the binary inclusion vector $\mathbf{b}\in\{0,1\}^{K}$ where $b_{k}=1$ if trajectory $\tau_{k}$ is included.
-The MPPI-Datamodel predicts weighted cost from sample membership via the linear model
-$g_{\theta}(\mathbf{b})=\theta^{\top}\mathbf{b}+\theta_{0}\approx\bar{C}(\mathcal{T}^{\prime})$ (13)
-where the coefficient $\theta_{k}$ represents the marginal influence of sample $\tau_{k}$ on the weighted cost.
-Since the MPPI control update $u^{*}=\sum_{k}w_{k}u_{k}$ uses the same importance weights $w_{k}$, samples that significantly affect the weighted cost also significantly affect the control output.
-This justifies using influence on weighted cost as a proxy for influence on the control decision.
+$\delta u_{0}(\mathcal{T}^{\prime})=\sum_{k:\tau_{k}\in\mathcal{T}^{\prime}}w_{k}(\mathcal{T}^{\prime})\,\epsilon_{0,k}$ (9)
+where $\epsilon_{0,k}$ is the first-step control perturbation of trajectory $\tau_{k}$.
+To capture the relationship between sample inclusion and the control update, we define the binary inclusion vector $\mathbf{b}\in\{0,1\}^{K}$ where $b_{k}=1$ if trajectory $\tau_{k}$ is included.
+The MPPI-Datamodel predicts the control update from sample membership via the linear model
+$g_{\theta}(\mathbf{b})=\theta^{\top}\mathbf{b}+\theta_{0}\approx\delta u_{0}(\mathcal{T}^{\prime})$ (10)
+where we fit separate regressions for each control dimension and average the absolute coefficients: $|\theta_{k}|=\frac{1}{m}\sum_{d=1}^{m}|\theta_{k}^{(d)}|$.
+The resulting $|\theta_{k}|$ quantifies the marginal influence of sample $\tau_{k}$ on the immediate control action.
+Samples with large $|\theta_{k}|$ disproportionately affect the first control applied to the system.
 III-B Offline Training Phase
@@
 We generate $M$ random subsets by including each trajectory independently with probability $\alpha\in(0,1)$.
-For each subset $j$ with inclusion mask $\mathbf{b}_{j}\in\{0,1\}^{K}$, we compute the subset-restricted weighted cost
-$\bar{C}_{j}^{(i)}=\frac{\sum_{k=1}^{K}b_{j,k}\,w_{k}^{(i)}\,C_{k}^{(i)}}{\sum_{k=1}^{K}b_{j,k}\,w_{k}^{(i)}}$ (14)
-where $w_{k}^{(i)}=\exp(-C_{k}^{(i)}/\lambda)$ are the unnormalized importance weights. The influence coefficients are obtained by solving the LASSO regression problem
-$\theta^{(i)*}=\operatorname*{arg\,min}_{\theta,\theta_{0}}\frac{1}{M}\sum_{j=1}^{M}\left(\theta^{\top}\mathbf{b}_{j}+\theta_{0}-\bar{C}_{j}^{(i)}\right)^{2}+\mu\|\theta\|_{1}$ (15)
-where $\theta_{0}$ is an intercept term excluded from regularization and the $\ell_{1}$ penalty encourages sparsity.
+For each subset $j$ with an inclusion mask $\mathbf{b}_{j}\in\{0,1\}^{K}$, we compute the subset-restricted control update $\delta u_{0,j}^{(i)}\in\mathbb{R}^{m}$ by renormalizing the MPPI weights over the included trajectories and taking the weighted sum of first-step perturbations.
+The influence coefficients for each control dimension $d$ are obtained by solving
+$\theta^{(i,d)*}=\operatorname*{arg\,min}_{\theta,\theta_{0}}\frac{1}{M}\sum_{j=1}^{M}\left(\theta^{\top}\mathbf{b}_{j}+\theta_{0}-[\delta u_{0,j}^{(i)}]_{d}\right)^{2}+\mu\|\theta\|_{1}$ (11)
+where $\theta_{0}$ is an intercept term excluded from regularization, and the $\ell_{1}$ penalty encourages sparsity.
+We then compute a single influence score per sample by averaging absolute coefficients across dimensions: $|\theta_{k}^{(i)*}|=\frac{1}{m}\sum_{d=1}^{m}|\theta_{k}^{(i,d)*}|$.
 The key to recovering individual influences is the randomness of the subsets. For any two trajectories $k\neq k^{\prime}$, their inclusion events are independent across subsets:
 $\mathbb{E}[b_{j,k}\cdot b_{j,k^{\prime}}]=\mathbb{E}[b_{j,k}]\cdot\mathbb{E}[b_{j,k^{\prime}}]=\alpha^{2}$ (12)
-This independence decorrelates the contributions of different samples, allowing the regression to disentangle individual effects from aggregate subset effects.
-With sufficient subsets $M$, the LASSO solution $\theta_{k}^{(i)*}$ accurately estimates the marginal influence of each sample $k$.
+This independence decorrelates the contributions of different samples, helping the regression disentangle individual effects from aggregate subset effects.
+The LASSO solution $\theta_{k}^{(i)*}$ provides an approximate influence score for each sample $k$; the quality of this approximation depends on $M$ and on the degree of nonlinearity in the control update mapping (Section II-C).
 Step 3: Influence Predictor Training.
@@
 Intuitively, a trajectory with cost $C_{k}=10$ has high influence if all other trajectories have costs above $100$, but low influence if other trajectories have similar costs around $10$.
-To capture this context dependence, we augment the input features with summary statistics:
-$\hat{\theta}_{k}=h_{\phi}\left(C_{k},C_{\text{viol},k},\bar{C}^{(i)},\sigma_{C}^{(i)}\right)$ (17)
-where $\bar{C}^{(i)}=\frac{1}{K}\sum_{k=1}^{K}C_{k}^{(i)}$ is the mean cost and $\sigma_{C}^{(i)}=\text{std}(\{C_{k}^{(i)}\}_{k=1}^{K})$ is the standard deviation.
-These statistics provide context about where sample $k$ sits relative to other samples in the instance.
-The predictor parameters are learned by minimizing prediction error across all collected instances:
-$\phi^{*}=\operatorname*{arg\,min}_{\phi}\sum_{i=1}^{N}\sum_{k=1}^{K}\left(h_{\phi}\left(C_{k}^{(i)},C_{\text{viol},k}^{(i)},\bar{C}^{(i)},\sigma_{C}^{(i)}\right)-\theta_{k}^{(i)*}\right)^{2}$ (18)
+To capture this context dependence, we construct an 8-dimensional feature vector:
+$\hat{\theta}_{k}=h_{\phi}\left(\mathbf{f}_{k}^{(i)}\right)$ (13)
+where $\mathbf{f}_{k}^{(i)}=[C_{k},\;C_{\text{viol},k},\;z_{k},\;\tilde{C}_{k},\;s_{k},\;\mathbf{1}_{C_{\text{viol},k}>0},\;\bar{C}^{(i)},\;\sigma_{C}^{(i)}]^{\top}$.
+Here $z_{k}=(C_{k}-\bar{C})/\sigma_{C}$ is the $z$-score, $\tilde{C}_{k}=(C_{k}-C_{\min})/\sigma_{C}$ is the normalized distance from the best sample, $s_{k}=\exp(-(C_{k}-C_{\min})/\sigma_{C})$ is a soft exponential rank, $\bar{C}^{(i)}=\frac{1}{K}\sum_{k}C_{k}^{(i)}$ is the mean cost, and $\sigma_{C}^{(i)}=\text{std}(\{C_{k}^{(i)}\})$ is the cost standard deviation.
+The per-sample features ($z_{k},\tilde{C}_{k},s_{k}$) encode where sample $k$ sits in the cost distribution, while the instance statistics ($\bar{C},\sigma_{C}$) provide context.
+The predictor parameters are learned by minimizing the prediction error across all collected instances:
+$\phi^{*}=\operatorname*{arg\,min}_{\phi}\sum_{i=1}^{N}\sum_{k=1}^{K}\left(h_{\phi}\left(\mathbf{f}_{k}^{(i)}\right)-|\theta_{k}^{(i)*}|\right)^{2}$ (14)
 In practice, $h_{\phi}$ can be a small neural network (e.g., two hidden layers with 64 units) or even a linear model if the relationship is sufficiently simple.
@@
 9:   Sample inclusion mask: $b_{j,k}\sim\text{Bernoulli}(\alpha)$ for all $k$
-10:   Compute weighted cost: $\bar{C}_{j}^{(i)}\leftarrow\frac{\sum_{k}b_{j,k}\,w_{k}^{(i)}\,C_{k}^{(i)}}{\sum_{k}b_{j,k}\,w_{k}^{(i)}}$
-11:  end for
-12:  Solve LASSO (15) to obtain $\theta^{(i)*}\in\mathbb{R}^{K}$
-13:  for $k=1$ to $K$ do
-14:   $\mathcal{D}\leftarrow\mathcal{D}\cup\{(C_{k}^{(i)},C_{\text{viol},k}^{(i)},\bar{C}^{(i)},\sigma_{C}^{(i)},\theta_{k}^{(i)*})\}$
-15:  end for
-16:end for
-17: Train $h_{\phi^{*}}$ on $\mathcal{D}$ by minimizing (18)
-18:return $h_{\phi^{*}}$
+10:   Compute subset control update: $\delta u_{0,j}^{(i)}\leftarrow\sum_{k:b_{j,k}=1}\tilde{w}_{k}^{(i,j)}\,\epsilon_{0,k}^{(i)}$
+11:    with $\tilde{w}_{k}^{(i,j)}=\exp(-C_{k}^{(i)}/\lambda)/\sum_{l:b_{j,l}=1}\exp(-C_{l}^{(i)}/\lambda)$
+12:  end for
+13:  Solve LASSO (11) per control dimension, average $|\theta_{k}^{(i)*}|=\frac{1}{m}\sum_{d}|\theta_{k}^{(i,d)*}|$
+14:  for $k=1$ to $K$ do
+15:   Build feature vector $\mathbf{f}_{k}^{(i)}$ per (13)
+16:   $\mathcal{D}\leftarrow\mathcal{D}\cup\{(\mathbf{f}_{k}^{(i)},|\theta_{k}^{(i)*}|)\}$
+17:  end for
+18:end for
+19: Train $h_{\phi^{*}}$ on $\mathcal{D}$ by minimizing (14)
+20:return $h_{\phi^{*}}$
 III-C Online Inference Phase
@@
 Importantly, the actual MPPI control computation still operates on the total cost with the current penalty $\rho$, which can be adjusted online without retraining the predictor.
-Given a new set of $K$ sampled trajectories, we first compute the cost components $(C_{\text{goal},k},C_{\text{ctrl},k},C_{\text{viol},k})$ and total costs $C_{k}=C_{\text{goal},k}+C_{\text{ctrl},k}+\rho\cdot C_{\text{viol},k}$.
-We then compute the instance statistics
-$\bar{C}=\frac{1}{K}\sum_{k=1}^{K}C_{k},\qquad\sigma_{C}=\text{std}(\{C_{k}\}_{k=1}^{K})$ (19)
-and obtain the predicted influence coefficients as
-$\hat{\theta}_{k}=h_{\phi^{*}}(C_{k},C_{\text{viol},k},\bar{C},\sigma_{C})$ (20)
+Given a new set of $K$ sampled trajectories, we compute cost components and build the 8-dimensional feature vector $\mathbf{f}_{k}$ for each sample as in (13). The predicted influence magnitudes are
+$\hat{\theta}_{k}=h_{\phi^{*}}(\mathbf{f}_{k})$ (15)
 These predictions identify which samples contribute meaningfully to the control output.
-Sample Pruning for Efficiency. We retain only those samples whose predicted influence magnitude exceeds a threshold $\tau$:
-$\mathcal{K}_{\text{keep}}=\{k:|\hat{\theta}_{k}|\geq\tau\}$ (21)
-This pruning step reduces computational cost while preserving the samples that matter most. The MPPI control update is then computed using only the retained samples:
+Sample Pruning for Efficiency. We retain the top $n_{\text{keep}}$ samples ranked by predicted influence magnitude:
+$\mathcal{K}_{\text{keep}}=\text{top-}n_{\text{keep}}\{k:|\hat{\theta}_{k}|\}$ (16)
+This fixed-size pruning step preserves the samples that matter most for the control update. The MPPI control update is then computed using only the retained samples:
 $u^{*}=\sum_{k\in\mathcal{K}_{\text{keep}}}\tilde{w}_{k}\,u_{k},\qquad\tilde{w}_{k}=\frac{\exp(-C_{k}/\lambda)}{\sum_{j\in\mathcal{K}_{\text{keep}}}\exp(-C_{j}/\lambda)}$ (17)
 where the importance weights are renormalized over $\mathcal{K}_{\text{keep}}$.
-Adaptive Constraint Penalty for Safety. To adapt the constraint penalty online, we monitor the influence contributed by constraint-violating samples. We define the violation influence ratio as
-$r_{\text{viol}}=\frac{\sum_{k\in\mathcal{K}_{\text{viol}}}|\hat{\theta}_{k}|}{\sum_{k=1}^{K}|\hat{\theta}_{k}|}$ (23)
-where $\mathcal{K}_{\text{viol}}=\{k:C_{\text{viol},k}>0\}$ denotes the set of constraint-violating samples.
-This ratio quantifies how much of the total influence comes from samples that violate constraints.
-When $r_{\text{viol}}$ is high, constraint-violating trajectories disproportionately affect the control output, suggesting the current penalty $\rho$ is insufficient to discourage unsafe behaviors.
-We update the penalty coefficient according to
-$\rho_{t+1}=\rho_{t}+\eta\cdot(r_{\text{viol}}-r_{\text{target}})$ (24)
-where $\eta>0$ is a step size and $r_{\text{target}}\in(0,1)$ is the desired maximum violation influence ratio.
-This adaptation increases $\rho$ when violation influence exceeds the target and decreases it otherwise, maintaining constraint satisfaction while avoiding overly conservative behavior.
-Algorithm 2 presents the complete online procedure.
+Adaptive Constraint Penalty for Safety. To adapt the constraint penalty online, we monitor the fraction of sampled trajectories that violate constraints. We define the violation rate as
+$r_{\text{viol}}=\frac{1}{K}\sum_{k=1}^{K}\mathbb{1}[C_{\text{viol},k}>0]$ (18)
+and maintain an exponential moving average $\hat{r}_{\text{viol},t}=(1-\alpha_{\text{ema}})\hat{r}_{\text{viol},t-1}+\alpha_{\text{ema}}\,r_{\text{viol},t}$ with $\alpha_{\text{ema}}=0.1$.
+When $\hat{r}_{\text{viol}}$ exceeds the target, many trajectories violate constraints, suggesting the penalty $\rho$ is insufficient.
+We update
+$\rho_{t+1}=\text{clip}\left(\rho_{t}+\eta\cdot(\hat{r}_{\text{viol},t}-r_{\text{target}}),\;\rho_{\min},\;\rho_{\max}\right)$ (19)
+where $\eta>0$ is a step size and $r_{\text{target}}\in(0,1)$ is the desired maximum violation rate. Algorithm 2 presents the complete online procedure.
 Algorithm 2 Online: Datamodel-Based MPPI
-0: Trained predictor $h_{\phi^{*}}$, initial state $x_{0}$, nominal control $\bar{U}$, initial penalty $\rho$, pruning threshold $\tau$, target violation ratio $r_{\text{target}}$, adaptation step size $\eta$, temperature $\lambda$
-1:for each control iteration do
-2:  Sample $K$ trajectories from current state
-3:  Compute cost components: $(C_{\text{goal},k},C_{\text{ctrl},k},C_{\text{viol},k})$ for all $k$
-4:  Compute total costs: $C_{k}\leftarrow C_{\text{goal},k}+C_{\text{ctrl},k}+\rho\cdot C_{\text{viol},k}$
-5:  Compute instance statistics: $\bar{C}\leftarrow\frac{1}{K}\sum_{k}C_{k}$,  $\sigma_{C}\leftarrow\text{std}(\{C_{k}\})$
-6:  Predict influence: $\hat{\theta}_{k}\leftarrow h_{\phi^{*}}(C_{k},C_{\text{viol},k},\bar{C},\sigma_{C})$ for all $k$
-7:  Identify kept samples: $\mathcal{K}_{\text{keep}}\leftarrow\{k:|\hat{\theta}_{k}|\geq\tau\}$
-8:  Identify violating samples: $\mathcal{K}_{\text{viol}}\leftarrow\{k:C_{\text{viol},k}>0\}$
-9:  Compute violation ratio: $r_{\text{viol}}\leftarrow\frac{\sum_{k\in\mathcal{K}_{\text{viol}}}|\hat{\theta}_{k}|}{\sum_{k=1}^{K}|\hat{\theta}_{k}|}$
-10:  Update penalty: $\rho\leftarrow\rho+\eta\cdot(r_{\text{viol}}-r_{\text{target}})$
-11:  Compute weights: $\tilde{w}_{k}\leftarrow\frac{\exp(-C_{k}/\lambda)}{\sum_{j\in\mathcal{K}_{\text{keep}}}\exp(-C_{j}/\lambda)}$ for $k\in\mathcal{K}_{\text{keep}}$
-12:  Compute optimal control: $u^{*}\leftarrow\sum_{k\in\mathcal{K}_{\text{keep}}}\tilde{w}_{k}\,u_{k}$
+0: Trained predictor $h_{\phi^{*}}$, initial state $x_{0}$, nominal control $\bar{U}$, initial penalty $\rho$, number to keep $n_{\text{keep}}$, target violation rate $r_{\text{target}}$, step size $\eta$, EMA parameter $\alpha_{\text{ema}}$, temperature $\lambda$
+[…]
+2:for each control iteration do
+3:  Sample $K$ perturbations $\{\epsilon_{k}\}$, roll out trajectories from current state
+4:  Compute cost components: $(C_{\text{goal},k},C_{\text{ctrl},k},C_{\text{viol},k})$ for all $k$
+5:  Compute total costs: $C_{k}\leftarrow C_{\text{goal},k}+C_{\text{ctrl},k}+\rho\cdot C_{\text{viol},k}$
+6:  Build features $\mathbf{f}_{k}$ and predict influence: $\hat{\theta}_{k}\leftarrow h_{\phi^{*}}(\mathbf{f}_{k})$
+7:  Select top $n_{\text{keep}}$ samples by $|\hat{\theta}_{k}|$: $\mathcal{K}_{\text{keep}}$
+8:  Compute violation rate: $r_{\text{viol}}\leftarrow\frac{1}{K}\sum_{k}\mathbf{1}[C_{\text{viol},k}>0]$
+9:  Update EMA: $\hat{r}_{\text{viol}}\leftarrow(1-\alpha_{\text{ema}})\hat{r}_{\text{viol}}+\alpha_{\text{ema}}\,r_{\text{viol}}$
+10:  Update penalty: $\rho\leftarrow\text{clip}(\rho+\eta(\hat{r}_{\text{viol}}-r_{\text{target}}),\rho_{\min},\rho_{\max})$
+11:  Recompute weights over $\mathcal{K}_{\text{keep}}$: $\tilde{w}_{k}\leftarrow\frac{\exp(-C_{k}/\lambda)}{\sum_{j\in\mathcal{K}_{\text{keep}}}\exp(-C_{j}/\lambda)}$
+12:  Compute control: $u^{*}\leftarrow\bar{u}+\sum_{k\in\mathcal{K}_{\text{keep}}}\tilde{w}_{k}\,\epsilon_{k}$
 13:  Execute first control $u_{0}^{*}$, shift nominal sequence $\bar{U}$
 14:end for
-Remark 1 (Unified Influence for Efficiency and Safety).
-Because the predictor takes both total cost and violation cost as input, a single learned model captures influence relevant to both sample efficiency and constraint handling.
-The pruning mechanism improves efficiency by discarding low-influence samples regardless of their constraint status, while the adaptive $\rho$ mechanism improves safety by strengthening penalties when constraint-violating samples remain influential.
-These two mechanisms operate complementarily: pruning reduces computation without sacrificing control quality, while penalty adaptation ensures the kept samples lead to safe behaviors.
-Remark 2 (Computational Complexity).
-The offline phase requires $O(N\cdot M\cdot K)$ subset evaluations and $N$ LASSO solves, but is performed only once.
-The online phase requires only $O(K)$ forward passes through $h_{\phi^{*}}$ per control iteration, making influence prediction negligible compared to trajectory sampling and cost evaluation.
-The pruning step further reduces the cost of the weighted average computation from $O(K)$ to $O(|\mathcal{K}_{\text{keep}}|)$.
+Remark 1 (Pruning and Penalty Adaptation)
+The pruning step and the penalty adaptation work independently of each other.
+Pruning keeps the samples with the highest predicted influence on the control update, so the weighted average is computed over fewer but more informative samples.
+The adaptive penalty tracks how often sampled trajectories violate constraints and adjusts $\rho$ to compensate.
+It is worth noting that pruning by influence magnitude does not automatically favor safe trajectories — a sample with high influence can still violate constraints (see Section IV).
+Remark 2 (Computational Complexity)
+The offline phase involves $O(N\cdot M\cdot K)$ subset evaluations and $N$ LASSO solves, but it only needs to be done once.
+At runtime, the online phase only requires $O(K)$ forward passes through $h_{\phi^{*}}$ per control step, which is small compared to the cost of sampling trajectories and evaluating costs.
+Pruning also reduces the weighted average computation from $O(K)$ down to $O(|\mathcal{K}_{\text{keep}}|)$.
 IV Experiments
-We evaluate the proposed framework on a path tracking task with obstacle avoidance, demonstrating unified benefits for sample efficiency and constraint-aware control.
-IV-A Setup
-We consider a vehicle governed by the kinematic bicycle model with state $x=[p_{x},p_{y},\psi,v]^{\top}$ (position, heading, velocity) and control $u=[\delta,a]^{\top}$ (steering angle, acceleration).
-The vehicle tracks an oval reference path while avoiding circular obstacles.
-The goal cost penalizes deviation from the reference path, while the violation cost applies a quadratic penalty when the vehicle enters the obstacle regions.
-The MPPI controller uses a planning horizon of $T=20$ steps at $dt=0.1$ s and temperature $\lambda=100$.
-Control noise is sampled from $\mathcal{N}(0,\Sigma)$ with $\Sigma=\text{diag}(0.1,0.5)$ for steering and acceleration respectively.
-For the influence predictor $h_{\phi}$, we use a multilayer perceptron (MLP) with two hidden layers of 64 units each and ReLU activations.
-The network takes as input the normalized costs $(S_{k},S_{\text{viol},k},\bar{S},\sigma_{S})$ and outputs the predicted influence coefficient $\hat{\theta}_{k}$.
-Training uses the Adam optimizer with learning rate $10^{-3}$ for 1000 epochs.
-The datamodel parameters are set as: inclusion probability $\alpha=0.5$, number of subsets $M=50$, and LASSO regularization $\mu=0.01$.
-We collect $N=200$ MPPI instances from randomly sampled initial states for offline training.
-IV-B Sample Efficiency Analysis
-Figure 2: Sample efficiency comparison between DM-MPPI and standard MPPI across varying sample sizes $K\in\{50,100,150,\ldots,500\}$.
-To determine the minimum number of samples required for effective control, we evaluate both methods across sample sizes $K\in\{50,100,150,\ldots,500\}$.
-Figure 2 presents the results.
-Standard MPPI shows a steep performance degradation below $K=200$, with position RMSE increasing from 0.41 m at $K=500$ to 1.30 m at $K=50$.
-In contrast, DM-MPPI maintains stable performance across a wider range: at $K=100$, it achieves position RMSE of 0.44 m and $r_{\text{viol}}=0.11$, comparable to standard MPPI at $K=500$ (0.41 m and 0.09 respectively).
-However, the datamodel approach has a minimum sample requirement.
-At $K=50$, DM-MPPI performance degrades noticeably (position RMSE increases to 0.72 m, $r_{\text{viol}}$ rises to 0.35), suggesting insufficient samples for reliable influence estimation.
-This aligns with the effective influence size $K_{\text{eff}}^{\theta}\approx 28$ observed in our analysis—when the total sample count approaches this threshold, the datamodel cannot effectively distinguish high-influence samples from noise.
-Based on this analysis, we select $K=100$ for DM-MPPI in subsequent experiments, achieving $5\times$ sample reduction compared to standard MPPI ($K=500$) while maintaining equivalent control quality.
-IV-C Performance Comparison
-Figure 3: Comparison of sampled different trajectories.
-TABLE I: Performance Comparison
-Metric Standard DM-Fixed DM-Adaptive
-Samples used 500 100 100
-Position RMSE (m) 0.41 0.44 0.43
-Heading RMSE (rad) 0.031 0.033 0.032
-Min. obstacle dist. (m) 1.24 1.31 1.42
-$r_{\text{viol}}$ 0.091 0.082 0.071
-Iteration time (ms) 85.4 18.2 18.5
-We compare three configurations: (1) Standard MPPI with $K=500$ samples and fixed penalty $\rho=10^{10}$, (2) DM-MPPI with $K=100$ samples and fixed $\rho=10^{10}$, and (3) DM-MPPI with $K=100$ samples and adaptive $\rho$ using $r_{\text{target}}=0.05$ and step size $\eta=10^{9}$.
-Table I summarizes the performance comparison.
-The datamodel approach enables $5\times$ sample reduction while maintaining tracking quality.
-Figure 3 visualizes the sampled trajectories at a representative timestep; despite using only 100 samples, the datamodel variant produces a control output comparable to standard MPPI with 500 samples.
-Analysis of the learned influence coefficients reveals an effective influence size of $K_{\text{eff}}^{\theta}\approx 28$, confirming that roughly one-quarter of the samples carry meaningful influence while the remainder can be safely pruned.
-The adaptive variant achieves the largest minimum obstacle distance by dynamically adjusting $\rho$ based on the violation influence ratio $r_{\text{viol}}$.
-V Limitations and Future Work
-Soft Constraint Guarantees.
-The proposed framework relies on soft constraint penalties, which cannot guarantee hard constraint satisfaction.
-While the adaptive $\rho$ mechanism reduces the influence of constraint-violating samples, it cannot ensure that the resulting control will never violate constraints.
-Safety-critical applications may require integration with formal methods such as Control Barrier Functions [20, 21] or tube-based robust MPC [17].
-Future work could integrate the datamodel framework with CBFs, using predicted influence to inform barrier parameters.
-Limited Experimental Scope.
-The current evaluation considers a single path tracking task with obstacle avoidance using a kinematic bicycle model.
-While this serves as a representative benchmark, comprehensive validation requires evaluation across diverse domains including manipulation, legged locomotion, and higher-dimensional systems with more complex constraint structures.
-Future work will include comparisons against state-of-the-art sample-efficient MPPI variants [7, 16], evaluation under varying noise levels and constraint complexities, and deployment on real hardware platforms.
-Fixed Predictor After Offline Training.
-The influence predictor is fixed after offline training, which limits adaptability to changing conditions.
-Online fine-tuning would allow the predictor to adapt to distribution shifts or novel scenarios, but introduces computational overhead that conflicts with real-time requirements.
-One promising direction is to maintain a small replay buffer of recent MPPI instances and perform incremental updates during idle cycles.
-Alternatively, meta-learning approaches could train the predictor to rapidly adapt to new conditions with minimal fine-tuning samples.
-Generalization Across Cost Scales.
-The framework uses a fixed inclusion probability $\alpha$ during offline training, and the learned predictor generalizes well within the range of cost distributions encountered during training.
-However, performance may degrade when deployed in scenarios where cost statistics differ substantially from training.
-Normalizing costs by instance-level statistics partially addresses this issue, but does not fully resolve it when the underlying cost structure changes significantly.
-Future work could explore scale-invariant predictor architectures that operate on relative cost rankings rather than absolute values to enable broader generalization.
-VI Conclusion
-We presented a framework that extends the datamodeling paradigm from machine learning to sampling-based optimal control.
-While traditional datamodels estimate sample influence through regression on a fixed dataset, our approach trains an influence predictor that generalizes across control instances—enabling real-time influence estimation without solving regression problems online.
-By learning to map cost features to influence coefficients, the predictor identifies which samples contribute meaningfully to the control output, allowing efficient sample pruning and adaptive constraint handling through a single learned model.
-The key insight is that influence coefficients, though computed via subset regression during offline training, show predictable structure as a function of sample costs and global statistics.
-This structure enables amortized inference: rather than re-running regression for each new MPPI instance, we predict influence directly from cost features in a single forward pass.
-Experiments on path tracking with obstacle avoidance demonstrate $5\times$ sample reduction while maintaining tracking performance and achieving improved safety margins through adaptive penalty tuning.
+We evaluate the proposed framework on two closed-loop tracks: a heart-shaped course and a figure-8 course, with static obstacle avoidance using a kinematic bicycle model.
+All methods share identical dynamics and track geometry; they differ only in sample count and cost/sampling strategy.
+All experiments use 8 random seeds and report the mean $\pm$ standard deviation.
+IV-A Experimental Setup
+Dynamics. We use a kinematic bicycle model with state $x=[p_{x},p_{y},\psi,v]^{\top}$ and control $u=[a,\delta]^{\top}$ (acceleration, steering angle):
+$p_{x,t+1}$ $=p_{x,t}+v_{t}\cos(\psi_{t})\,\Delta t$
+$p_{y,t+1}$ $=p_{y,t}+v_{t}\sin(\psi_{t})\,\Delta t$
+$\psi_{t+1}$ $=\psi_{t}+\frac{v_{t}}{L}\tan(\delta_{t})\,\Delta t$
+$v_{t+1}$ $=\text{clip}(v_{t}+a_{t}\,\Delta t,\;v_{\min},\;v_{\max})$ (20)
+with wheelbase $L=0.3$ m, $\Delta t=0.1$ s, $v_{\max}=2.5$ m/s, $v_{\min}=0.3$ m/s, and maximum steering angle $|\delta|\leq 0.55$ rad.
+Track environments. We consider two closed-loop reference paths:
+Heart track. A modified parametric heart curve:
+$h_{x}(t)$ $=s\cdot 16\sin^{3}(t),$
+$h_{y}(t)$ $=s\cdot(13\cos t-3\cos 2t-0.8\cos 3t-0.2\cos 4t)$ (21)
+with scale $s=0.6$, centered at $(5,5)$ m.
+The track has an arc length of $53.3$ m, a constant width of $0.7$ m, and ten rectangular obstacles ($0.40\times 0.28$ m) placed at evenly spaced arc length fractions with alternating lateral offsets.
+Lemniscate track. A lemniscate of Bernoulli:
+$\ell_{x}(t)$ $=\frac{a\cos t}{1+\sin^{2}t},\qquad\ell_{y}(t)=\frac{a\sin t\cos t}{1+\sin^{2}t}$ (22)
+with $a=4.5$, centered at $(5,5)$ m.
+The figure-8 has a self-intersection where the track narrows to $30\%$ of the base width ($0.7$ m).
+Eight obstacles are placed along the path with tighter lateral offsets near the crossing.
+IV-B Baselines
+We compare nine methods spanning standard MPPI at multiple sample budgets, heuristic pruning, and the proposed datamodel-based variants (Fig.
+2).
+MPPI K=50/100/200/500K\!=\!50/100/200/500 runs standard MPPI with i.i.d.
+samples; K=50K\!=\!50 uses only quadratic centerline cost ($w_{\text{track}}=8.0$), obstacle barriers, and a progress reward, while $K\geq 100$ adds boundary violation and heading alignment penalties.
+All four pruning baselines start with K=500K\!=\!500 samples and keep 100.
+Random Pruning simply picks 100 at random, giving a lower bound on what pruning can do.
+Top-Weight Pruning keeps the 100 samples that already carry the most MPPI importance weight $w_{k}$—in other words, the trajectories that would dominate the weighted average anyway.
+Cost-Threshold Pruning throws out the 400 most expensive trajectories and recomputes MPPI weights over the remaining 100.
+Elite-CEM also keeps the 100 cheapest samples, but instead of applying MPPI weights, it just averages their perturbations uniformly, following the Cross-Entropy Method idea of working with an elite set.
+DM-fixed builds on MPPI K=500K\!=\!500 with temporally correlated AR(1) noise ($\beta=0.6$), 3 MPPI iterations per timestep, control smoothness penalty, curvature-aware target speed, and influence-based pruning.
+DM-adapted extends DM-fixed with adaptive temperature scaling, log-barrier boundary cost, look-ahead curvature-aware speed, ESS-driven noise adaptation, and elite-sample refinement.
+Figure 2: Representative trajectories on the heart track (left) and lemniscate track (right).
+IV-C Results
+TABLE I: Results across both tracks (8 seeds). Pruning methods draw K=500K\!=\!500 and retain 100 samples.
+Method Tracking RMSE Heading RMSE Boundary
+(m) (rad) Viol.
+Heart track (10 obstacles, arc length 53.3 m)
+MPPI K=50K\!=\!50 $0.65\pm 0.10$ $0.39\pm 0.12$ $88.7\pm 12.1$
+MPPI K=100K\!=\!100 $8\pm 0.06 0.33±0$$3\pm 0.08$ $31.3\pm 8.5$
+MPPI K=200K\!=\!200 $0.25\pm 0.04$ $0.30\pm 0.07$ $6.0\pm 5.2$
+MPPI K=500K\!=\!500 $0.18\pm 0.02$ $0.28\pm 0.01$ $0.7\pm 0.6$
+Random Prune $0.71\pm 0.06$ $0.42\pm 0.08$ $97.3\pm 14.8$
+Top-Weight $0.21\pm 0.03$ $0.26\pm 0.10$ $2.0\pm 1.7$
+Cost-Threshold $0.24\pm 0.04$ $0.26\pm 0.08$ $6.0\pm 2.7$
+Elite-CEM $0.28\pm 0.06$ $0.26\pm 0.05$ $8.7\pm 1.5$
+DM-fixed $0.17\pm 0.02$ $\mathbf{0.23\pm 0.03}$ $1.2\pm 0.4$}
+DM-adapted $\mathbf{0.17\pm 0.01}$ $0.24\pm 0.05$ $\mathbf{0.8\pm 0.3}$
+Lemniscate track (8 obstacles, self-intersection)
+MPPI K=50K\!=\!50 $0.58\pm 0.10$ $0.36\pm 0.03$ $62.3\pm 10.4$
+MPPI K=100K\!=\!100 $0.34\pm 0.03$ $0.30\pm 0.02$ $22.7\pm 6.8$
+MPPI K=200K\!=\!200 $0.22\pm 0.02$ $0.27\pm 0.02$ $5.3\pm 4.1$
+MPPI K=500K\!=\!500 $0.16\pm 0.01$ $0.25\pm 0.01$ $1.3\pm 0.5$
+Random Prune $0.64\pm 0.07$ $0.39\pm 0.14$ $71.0\pm 11.6$
+Top-Weight $0.25\pm 0.07$ $0.23\pm 0.07$ $8.3\pm 1.2$
+Cost-Threshold $0.31\pm 0.05$ $0.23\pm 0.12$ $9.4\pm 4.2$
+Elite-CEM $0.39\pm 0.08$ $0.24\pm 0.05$ $7.0\pm 3.0$
+DM-fixed $0.19\pm 0.03$ $\mathbf{0.21\pm 0.08}$ $2.1\pm 0.7$}
+DM-adapted $\mathbf{0.16\pm 0.02}$ $0.22\pm 0.05$ $\mathbf{1.0\pm 0.5}$
+Table I presents results across both tracks.
+The clearest takeaway is that how you prune matters far more than whether you prune at all.
+Random pruning actually performs worse than every standard MPPI configuration, including the smallest sample budget K=50K\!=\!50—throwing out samples indiscriminately hurts the weighted average by dropping useful trajectories while keeping noisy ones.
+Informed pruning, on the other hand, brings clear improvements.
+Top-Weight, Cost-Threshold, and Elite-CEM all reduce tracking error compared to unpruned MPPI K=500K\!=\!500, though the gains vary across methods and tracks.
+All three achieve similar heading RMSE, which is expected since they are all selecting from the same pool of low-cost or high-weight samples.
+However, these heuristic methods still incur several boundary violations on both tracks, with the numbers getting worse on the lemniscate, where the narrowed self-intersection leaves less margin for error.
+DM-fixed and DM-adapted consistently achieve the fewest boundary violations across both tracks.
+DM-adapted reaches the best or near-best tracking RMSE on both courses while keeping violations roughly an order of magnitude lower than the heuristic pruning baselines.
+The correlated noise, multi-iteration refinement, and influence-based pruning in DM-fixed produce smoother trajectories that naturally stay further from boundaries, while the adaptive mechanisms in DM-adapted further sharpen tracking by slowing down early in tight sections.
+These trends hold across track geometries.
+The relative ordering of methods stays consistent on both tracks despite their different difficulties: the lemniscate’s self-intersection and variable width demand tighter control, yet the DM-adapted method keeps violations near one on average, while heuristic pruning methods see their violation counts climb.
+Finally, simply adding more samples shows diminishing returns—going from K=200K\!=\!200 to K=500K\!=\!500 barely changes tracking RMSE.
+Yet pruning K=500K\!=\!500 down to 100 well-chosen samples beats the full unpruned budget.
+A smaller set of good samples outperforms a larger set of unfiltered ones.
+V Conclusion
+We presented DM-MPPI, a framework that brings the datamodel idea from machine learning into MPPI control.
+We tested it on a kinematic bicycle model navigating two closed-loop tracks with obstacles.
+Informed sample pruning—keeping only the top $20\%$ of trajectories—substantially lowers tracking error compared to unpruned MPPI.
+Random pruning, by contrast, makes things worse than even the smallest sample budget.
+Top-weight, cost-threshold, and CEM-style pruning all improve tracking, but they still produce noticeably more boundary violations than the datamodel-based variants.
+DM-adapted consistently achieves the best or near-best tracking accuracy across both tracks while keeping boundary violations far below all other methods.
+This suggests that influence-aware selection offers a meaningful safety advantage over simple heuristic pruning.
+Future work includes: extending evaluation to higher-dimensional systems and dynamic obstacles to test generality; replacing the linear datamodel with nonlinear or attention-based influence predictors that capture sample interactions; developing online adaptation to mitigate training distribution mismatch; integrating with formal safety mechanisms such as Control Barrier Functions [1] or safe importance sampling [4] for hard constraint guaranties; coupling influence prediction with adaptive resampling, where the pruned budget is reallocated to resample in promising control regions; exploring how DM-MPPI can be combined with other MPPI improvements such as covariance adaptation and warm-starting.
+Acknowledgment
+Claude was used to assist with the language editing of this manuscript.